A theater contains 25 rows of seats. The first row contains 25 seats. Each succeeding row has one more seat than the row immediately before it. How many seats does the theater have?

Answer by richwmiller(17219)

25*(25)=625
12+25=300
625+300=925 seats
25+0
25+1
25+2
25+3
25+4
....
25+23
25+24
1 24
2 23
3 22